OpenStudy (anonymous):

Use synthetic division to find P(3) for: P(x)=3x^4-5x^2+6x-5

OpenStudy (anonymous):

@ganeshie8 do you know this? :/

ganeshie8 (ganeshie8):

uhh.. synthetic division

OpenStudy (anonymous):

yes im on my final for algebra II & i have no idea how to do this..

ganeshie8 (ganeshie8):

do u need to show work or u just need to tick the right option ?

OpenStudy (anonymous):

the right option

ganeshie8 (ganeshie8):

then we can cheat

ganeshie8 (ganeshie8):

P(x)=3x^4-5x^2+6x-5 P(3) = ?

ganeshie8 (ganeshie8):

just plugin x = 3 in that polynomial

ganeshie8 (ganeshie8):

it doesnt matter how u get the answer, as long as you get the same answer. both methods work and give the same answer.

OpenStudy (anonymous):

that sounds good to me ;o lol im almost graduated!

OpenStudy (anonymous):

wait what do you mean plug in x=3? im sorry

ganeshie8 (ganeshie8):

P(x)=3x^4-5x^2+6x-5 P(3) = 3(3)^4 - 5(3)^2 + 6(3) - 5 = ?

OpenStudy (anonymous):

lemme see

ganeshie8 (ganeshie8):

okie

OpenStudy (anonymous):

i got 211

ganeshie8 (ganeshie8):

correct !

ganeshie8 (ganeshie8):

and you can use wolfram to calculate things :-

OpenStudy (anonymous):

yay thank u again rex !!

ganeshie8 (ganeshie8):

welcome thanks for the cute testimonial :D

OpenStudy (anonymous):

no problem, u deserve it ^.^

ganeshie8 (ganeshie8):

ahh ty :)